ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 13, 10, 8 Colding, Tobias Holck, Courant Institute, New York University, USA Manifolds and spaces with Ricci curvature bounds

Results about manifolds and more generally spaces with Ricci curvature bounds will be presented. This will include geometric, topological and function theoretical results.

ICM 1998 Berlin · Plenary and Invited Lectures